A coin is released inside a lift at a height of 2m from the floor of the lift. The height of lift is 10m.the lift is moving with an acceleration of9m/s^2 downward.the time after which the coin will strike with lift is

Grade:12th pass

1 Answers

Shashi kant Singh
5 years ago
Net accln of coin is downwards 10-9=1m/s^2 so and distance covered to be 2m so s=ut+1/2at^2 hence 1m=0+1/2×10t^2hence t=root1/5